# Affordances and Design Principles of The Political Left and Right

### v1

@@ 3.3. Organisation of Co-design Workshops
The co-design workshop lasted about 2–3 hours and consisted of five activities that together (see Table 2). The first two activities were individual orientation and thinking activities. First, participants were asked to atune to the problems of social media and after, they were asked to experiment with the materiality of the social media service by considering its issues and challenges through primarly through an user interface, tasking users to consider its modifications. These activities were designed as individual to reduce group-think and peer pressure, allowing participants to form their own perspectives. Following the two individual activities, the participants were asked to share their thoughts and observations with the group and briefly introduce themselves if they had not done so already. This activity allowed all participants to share perspectives and observations without interruption. Following this, the participants moved on to the co-design activity, where they were assigned to design a new social media service in about 45–60 minutes. We outline the problem statement around political polarisation and decreasing political discussion culture, with the following brief:
@@
Presentation and survey Please present the prototype to the research team and answer their questions. Please fill in an exit survey. 15–25
@@
Your task is to work together to design a new service for people who want to discuss politics: PoliPulse. The service aims at politically active Finns to exchange opinions and reduce polarisation in society. The aim is to avoid the harmful features of other social media and to improve the functionality and atmosphere of social debate. You can only post text and links to the service. Your task as a group is to create an initial prototype of this service, showing the main interfaces and being able to explain what users do in these interfaces.
@@
In the final step, each group needed to present their joint proposal, which demanded that groups consolidate which ideas to include in their final artefact. We were primarily interested in these artefacts as the basis for our analysis, as they should exemplify ideological tendencies in the outcomes of a design process.
@@ 3.5. Analysis
The primary aim of this work was to highlight how political ideologies become embedded in the design process. The workshop we conducted served to gather the cultural norms, assumptions, and ideologies of the co-designers, here the political representatives, by capturing a design process. Co-design activities include examining various design directions and ideas; however, for this work, we wanted to focus on each group’s converged presentation. Therefore, we focus only on the co-designed artefacts, i.e. sketches on potential user interfaces (see Figure 3 for an example).
@@
We used qualitative content analysis to identify affordances from these sketches (Rodgers et al., 2000). We build on the Feature Analysis method (Hasinoff and Bivens, 2021) to connect this analysis to ideologies to examine design artefacts for their value-loaded contents. Hasinoff and Bivens (2021) developed the Feature Analysis as a qualitative content analysis that identifies ideologies in mobile applications based on their features. Their work builds on an inductive approach to identifying features (and later affordances) and grouping them to examine ideologies. As we already have ideological focuses of groups, we focus more on examining if there are patterns in the design approaches.
Naturally, the question becomes what is a feature or an affordance. For Hasinoff and Bivens (2021), a feature was “a function that users control or are likely aware of”. In contrast, an affordance was a higher level conceptualisation of a characteristic of the technology that “request, demand, allow, encourage, discourage, and refuse” (Davis and Chouinard, 2016, p. 241) a user behaviour. As affordances have been widely used in human–computer interaction and computer-mediated communication research (Evans et al., 2017; Hutchby, 2001), definitions vary. We built on scholars such as Boyd (2010), Bucher and Helmond (2017), DeVito et al. (2017), and Van Raemdonck and Pierson (2021), who have already identified various features and affordances on social media services, such as self-presentation (DeVito et al., 2017), communicative affordances (Bucher and Helmond, 2017), and persistence (Boyd, 2010). Mainly, we make use of the taxonomy by Van Raemdonck and Pierson (2021), on ”affordances that focus on norm building within the ‘networked publics’.” We use the taxonomy to categorise affordances along visibility, network interactions and intervention (Van Raemdonck and Pierson, 2021). In addition, we analysed which affordances structured the user’s self-presentation based on the framework by DeVito et al. (2017) under the category of profile work. After identifying features and categorising them into affordances, we compare affordances between groups. Using our understanding of how particular affordances describe normative relations to user behaviour, we consider them as value statements of the design groups. inline]More on how we actually did the coding work
@@ 4.1. General Architecture
Within forums, architectures varied. Group 1 structured the discussion forum through organisational units (such as health and welfare services, organised publicly and locally in Finland), followed by sub-forums based on ongoing projects (see Figure 3(b)). Group 2, on the other hand, structured the discussions into three groups: political parties, wider civic society, and other discussion. Similarly, Group 4 had a mixed approach based on thematic and location-based structures. With the structuring of the forum, Groups 1 and 4 added elements allowing citizen initiatives on establishing new sub-forums, showing a somewhat more citizen-centric approach to participation, going beyond just administratively defined forums.
The news feed-based design (see Figure 3(a)) did not have a preset structure but used hashtag-based taxonomy to allow user-generated structuring of the content.
@@ 4.2. Content and Posting
Only groups 3 and 4 specified how new content would be submitted to these forums. Both focused on more structured input opportunities to help citizens justify their positions. Group 4 designed an interface which scaffolds the post generation by asking user to enter their opinion and justifications on separated text input, with the assumption that this would lead to a higher quality debate (see Figure 4). In addition, they asked people to provide links to support their stance and recommended news content related to the opinion. Group 3 identified the same issue but considered an artificial intelligence agent (shown as a dog in Figure 3(a)) to suggest to the user that they could provide justifications. Their system also provided users with controls for who could view or react to their content. Interestingly, Group 3 introduced a character limit, focusing on shorter messages and propagation.
@@
Beyond providing written content and comments, two of the artefacts included a mechanism to react to the content. Group 1’s artefact used stars to show support, and such reactions would increase the topic’s visibility on lists based on what they called a “Reddit approach” This means there was competition for visibility, where more starred content would gain more visibility. Group 3 provided various reaction icons and the ability to recirculate content on their artefacts, with the aim of ‘allow showing support.’ This articulation is arguably different from Group 1’s, which did not consider reactions a mechanism for gaining popularity but emphasised their meaning for socio-emotional support in the final presentation. Groups 2 and 4 did not introduce this kind of affordance to the artefacts, with Group 2 articulating how reactions, such as ‘likes‘ are not beneficial for constructive discussion and lead to attention seeking.
Recent academic discussions on social media have focused on content moderation and its challenges (Rantasila et al., 2022). Remarkably, only groups 1 and 3 considered any form of content moderation in the final artefact, even though the focus was to devise an online space for political discussion. Group 1 hints at rules for the platform, which are enforced by moderators selected by some authority, which exemplifies the intention to govern the content (Rantasila et al., 2022). While it remains unclear who is part of the moderation team, the moderation must be conducted in a team effort. Group 3 describes a more user-driven approach to moderation with options to report posts, add trigger warnings, and show reminders to enforce rules and block users. Similarly, Group 3 mentions that there is a human editor to conduct fact-checking. This shows a mix of hard (e.g. removal of reported posts) and soft (e.g. trigger warnings) moderation styles (Rantasila et al., 2022). While we can surmise rules against spam and hate speech from Group 3 suggesting these as features, no group clearly described what content is warrant or sanctioned on their sites.
@@ 4.3. Profiles
All four designs contained the idea of a persistent user identity as a core part of the platform. At the same time, there was a lot of diversity in the details that should be included in the artefacts. All sites included some profile pictures. Group 2 allowed pseudonyms and a person to change their username, which would keep the posting history intact. Instead, groups 1, 3 and 4 enforced a real-name policy. However, Group 4 hinted at the possibility of parody accounts, although marked by an indicator. The policy of enforcing clear names is often debated also on social media services such as Facebook (Haimson and Hoffmann, 2016).
In terms of user verification, it was discussed in all final presentations. Group 1 argued that the system should have strong authentication as “the [discussion] data is used for political decision-making” but it also was an approach to ensure proper behaviour on the platform, and Group 3 argued in the same line that there should be a mechanism to ensure that the poster ‘is a real person’ and not pretend to be some other person. Similar arguments were also presented in Group 2 and Group 4.
Interestingly, the differences do not stop here (see Table 3). Profiles included a free-form description field in groups 3 and 4. Group 2 did not design such an element; instead, after the user name, an indication of the user’s party affiliation was shown. Group 4 added age and occupation in addition to the free-formed field, as well as a list of users’ political interests. Group 3 instead included, for example, the number of warnings the user had received for breaking any rules.
@@ 4.4. From Features to Affordances to Ideologies
The following section describes the differences and similarities of the artefacts along four key affordances (each row in Table 4). We identified the key affordances popularity-based visibility, open network interactions, user intervention and flexible profile work through categorising the features mentioned above based on current understanding of social media affordances (Boyd, 2010; Van Raemdonck and Pierson, 2021; DeVito et al., 2017). In the following paragraphs, we will go through each key affordance and the differences and commonalities between groups, along with their ideological leanings.
@@
Popularity-based Visibility Encourage Refuse Discourage
Open Network Interactions Refuse Allow Encourage Refuse
User Intervention Allow Encourage Encourage
@@
The groups varied in their stance towards the mechanics of content visibility on their platforms: both the right-leaning and mixed composition groups preferred popularity-based metrics to promote content on the platform. Posts would have a like count, prioritising their visibility on the platform. In Group 1, popular content was prominently displayed with stars on the front page. Therefore, both groups encouraged popularity-based visibility by fostering a usage pattern that aims for posts that receive more reactions and, therefore, get rewarded with more visibility, similar to platforms like Facebook (Van Raemdonck and Pierson, 2021). Both left-leaning Groups 2 and 3 rejected popularity-based design approaches. Strikingly, as we observed above, Group 3 mentions the need for reactions and likes, but they highlight them as easy ways to show support, not as a popularity-based logic. Therefore, Group 3 considered discouraging behaviour that rewards the popularity of content with higher visibility on the platform, whereas Group 2 outright refused it by not offering any way for users to engage with content based on its popularity.
@@
Regarding interaction forms, all but Group 3 opted for a subforum architecture. Group 1, 2 and 4 all decided to section the discussion on their platform according to organisational levels (Group 1); parties, society and others (Group 2); and themes or location (Group 4). Whether a platform technically allows anyone to see and interact with every user’s content, there is still a clear difference between a system requesting users to post in particular subforums and one accumulating posts on a news feed. The former can be considered closed group interactions because they happen in clearly delineated spaces, and content does not automatically reach users across a platform (Van Raemdonck and Pierson, 2021). Such architecture entrusts the site administrators and moderators responsible for sanctioning or rewarding certain behaviours. Contrary to that, the affordances of Group 3’s artefact enabled open network interactions, as content moves freely across the platform. Such an architecture focuses on the sender instead of an overarching group space like a subforum. This puts more responsibility and power onto the senders, as they have more possibilities for intervention and structuring the debate (Van Raemdonck and Pierson, 2021). This shows a contrast to the other left-leaning Group 2, as they chose the more conservative forum solution for their platform architecture.
@@
Regarding interventions, we observed various approaches and remarkable differences in the way the Groups’ artefact afforded users to be involved. The right-leaning Group 1 considered the moderation rules so important that one of their artefacts was a list of written rules. While their system imposed a governing style to moderation (Rantasila et al., 2022), they still allowed user interventions through the option to report to content and initiatives for new subforums, which shows a neutral stance. The left-leaning Group 3 instead used self-guided moderation strategies with trigger warnings and warning counts (which could lead to banning from the platform) (Rantasila et al., 2022). Further, together with the mixed Group 4, they thought of structured content input that promotes user in adding justification for their posts. Thus, both groups’ artefacts encourage intervention through users.
@@
Finally, it seems that groups including left-leaning participants provided more flexibility for users to present themselves, encouraging or at least not restricting flexible profile work (DeVito et al., 2017). While all groups required a persistent identity from their users, features such as a changeable username (Group 2) or parody accounts (Group 3) would encourage users to express themselves freely and differently. Additionally, Group 3 offered users control over the visibility of their content, encouraging users to perform self-presentation by catering to different audiences (DeVito et al., 2017). In stark contrast, the right-leaning group offered the least opportunities to customise profiles or present freely, which indicates an expectation for the users to present persistent roles on their platform and an affordance that discourages flexible profile work.
@@
Our study only had one right-leaning and one mixed composition group due to participant recruitment challenges (see foonote 1). However, two left-leaning groups already allow us to highlight diversity within the same end of the political spectrum, highlighting how various other factors may also impact design work. For example, with profile work, Group 3 was much more creative in terms of attributes (see Table 3). Further, Group 2 subscribes to the forum approach, which separates interactions on the platforms in pre-defined groups and gives less intervention capabilities to individual users.

### v2

@@ 3.3. Organisation of Co-design Workshops
The co-design workshop lasted about 2–3 hours and consisted of five activities altogether (see Table 2). The first two activities were individual orientation and thinking activities. First, participants were asked to attune towards the problems of social media and after they were asked to experiment with the materiality of the social media service by considering its issues and challenges primarily through a user interface, tasking users to consider its modifications. These activities were designed as individual ones to reduce group–think and peer pressure, allowing participants to form their own perspectives. Following the two individual activities, the participants were asked to share their thoughts and observations with the group and briefly introduce themselves if they had not done so already. This activity allowed all participants to share perspectives and observations without interruption. Following this, the participants moved on to the co-design activity, where they were assigned to design a new social media service in about 45–60 minutes. We outline the problem statement around political polarisation and decreasing political discussion culture, with the following brief:
@@
Presentation Please present the prototype to the research team and answer their questions. 15–25
@@
Your task is to work together to design a new service for people who want to discuss politics: PoliPulse. The service is aimed at politically active Finns to exchange opinions and reduce polarisation in society. The aim is to avoid the harmful features of other social media and to improve the functionality and atmosphere of social debate. You can only post text and links to the service. Your task as a group is to create an initial prototype of this service, showing the main interfaces and being able to explain what users do in these interfaces.
@@
In the final step, each group needed to present their joint proposal, which demanded that groups consolidate which ideas to include in their final artefact. We were primarily interested in these artefacts and their presentation as the basis for our analysis, as they should exemplify ideological tendencies in the outcomes of a design process.
In addition to the co-design activity, participants filled in a pre-survey (collecting basic demographics data) and an exit survey, asking about their experiences of the co-design workshop.
@@ 3.5. Analysis
The primary aim of this work was to highlight how political ideologies become embedded in the design process. The workshop we conducted served to gather the cultural norms, assumptions, and ideologies of the co-designers, here the political representatives, by capturing a design process. Co-design activities include examining various design directions and ideas; however, for this work, we wanted to focus on each group’s converged presentation. Therefore, we focus only on (1) the finished co-designed artefacts, i.e. sketches on potential user interfaces (see Figure 4 for an example) and (2) groups’ presentations on the artefacts and their operations in this analysis; i.e. the final step of the co-design workshops. For analysis purposes, the artefacts were scanned and final presentations were transcribed verbatim.
Artefact analysis
@@
We used qualitative content analysis to identify affordances from these sketches (Rodgers et al., 2000). We build on the Feature Analysis method (Hasinoff and Bivens, 2021) to connect this analysis to ideologies to examine design artefacts for their value-loaded contents. Hasinoff and Bivens (2021) developed the Feature Analysis as a qualitative content analysis that identifies ideologies in mobile applications based on their features. Their work builds on an inductive approach to identifying features and later categorising them into affordances.
Naturally, the question becomes what a feature or an affordance are and how we interpret them. For Hasinoff and Bivens (2021), a feature was “a function that users control or are likely aware of”. In contrast, for them an affordance was a higher level conceptualisation of a characteristic of the technology that “request, demand, allow, encourage, discourage, and refuse” (Davis and Chouinard, 2016, p. 241) a user behaviour.444 This list of verbs is an extension from the most classical notations of an affordance in human–computer interaction, where only two verbs might be used. We acknowledge that affordances have been widely used in human–computer interaction and computer-mediated communication research (Evans et al., 2017; Hutchby, 2001), with various definitions. Hasinoff and Bivens (2021) builds on Davis and Chouinard (2016) recent work on affordances, which has focused on a wider set of actions.
As such, an affordance describes the relationship between the system and the user on a more abstract level. As they promote or sanction specific user behaviours, they describe the norms set by the design. To examine affordances and their implicit norms specific to social media services, we examined works which classified affordances (Boyd, 2010; Bucher and Helmond, 2017; DeVito et al., 2017; Van Raemdonck and Pierson, 2021) to sensitise our analytical work on pre-existing abstracted affordances. These have included for example affordances related to self-presentation (DeVito et al., 2017), communicative functions (Bucher and Helmond, 2017), and persistence (Boyd, 2010). Two authors identified individual features through open coding of the material in multiple iterations of going over the material and discussing and refining codes following common qualitative coding practices. We then categorised those features into abstract affordances based on the existing literature. By examining the features and examining suitable higher-level abstractions, we ended up focusing on the affordances visibility, network interactions and intervention (Van Raemdonck and Pierson, 2021). In addition, we analysed which affordances structured the user’s self-presentation based on the framework by DeVito et al. (2017) under the category of profile work, a concept we adopted from (Silfverberg et al., 2011; Uski and Airi Lampinen, 2016) which refers to self-presentational labour done in social media settings in contrast to face-to-face situations.
Analysis of the final presentations
In addition to artefacts, we used a similar qualitative analysis process to examine the video recordings and verbatim transliterations of the presentation activity. This was the final step of the co-design workshop, where participants were asked to present the function of the service (using the artefacts) and the research team asked clarifying questions. We used this material to examine how groups motivated particular features and aligned our interpretation of the addition of said features (and related affordances) to how they described its intended use or reasoned its addition.
Mapping affordances and ideologies
Using our understanding of how particular affordances describe normative relations to user behaviour, we consider them as consolidated value statements of each group (see Figure 3). These value statements then align or contradict the behaviour in accordance to a political ideology.
Figure 3. Our analytical lens identified affordances and their implicit values to compare those to ideologies.
[…]
A diagram with text boxes and arrows pointing from left to right: Features enable affordances, affordances request, demand, allow, encourage, discourage, and refuse user behaviour which is encapsulated in a box stating value statement. This points towards to align or contradicting Ideologies.
Our final step was to map affordances into political ideologies. There is no established practices for this: Hasinoff and Bivens (2021) argues their research approach seeks to analyse ideologies, their analytical move from low-features and higher abstraction affordances (as described above) to ideologies is not clearly described. They suggest using speculative design approaches to map out alternatives from different ideological stances, however it remains unclear how ideologies are integrated into this analysis (for further highlights on this Section 5.3.)
Instead of seeking to interpret ideologies from the artifacts and material we contrasted the affordances across groups based on participants’ self-identified political ideology. In addition, we used existing literature from political science on the economic left–right and value conservative–liberal dimensions555 While we acknowledge there are various ways to define political ideologies (see Section 2.1), for this analytical stage we focused on these axes to clearly define what is meant by ideology in this work. Our interpretations on ideologies builld on previous work on these axis in Finland and Europe, such as those by Cochrane (2015a), Jahn (2023), Däubler and Benoit (2022) and Grönlund and Söderlund (2024). to establish the connection between choices on affordances and ideologies; or seek a connection with the identified affordance and political ideologies.. Naturally, there are other factors beyond party alignment (discussed in more detail in 5.3 and 5.4), this evidence provides justified interpretations of the role of political ideologies in the design of services; showing how we see a connection between an affordance, its potential description by participants, and political science work on ideologies. Therefore, as the final step we compared affordances between groups to identify patterns on affordances following political ideologies and connect them to previous work on ideologies.
@@ 4.1. General Architecture
[…]
A sketch showing a news feed designed by group 3
@@
Within forums, architectures varied. Group 1 structured the discussion forum through organisational units (such as health and welfare services, organised publicly and locally in Finland), followed by sub-forums based on ongoing projects (see Figure 4(b)). Group 2, on the other hand, structured the discussions into three groups: political parties, wider civic society, and other discussion. Similarly, Group 4 had a mixed approach based on thematic and location-based structures. With the structuring of the forum, Groups 1 and 4 added elements allowing citizen initiatives on establishing new sub-forums, showing a somewhat more citizen-centric approach to participation, going beyond just administratively defined forums. Group 3’s news feed-based design (see Figure 4(a)) did not have a preset structure but used hashtag-based taxonomy to allow user-generated structuring of the content.
@@ 4.2. Content and Posting
Only groups 3 and 4 specified how new content would be submitted to these forums. Both focused on more structured input opportunities to help citizens justify their positions. Group 4 designed an interface which scaffolds the post generation by asking the user to enter their opinion and justifications on separated text input, with the assumption that this would lead to a higher quality debate (see Figure 5). In addition, they asked people to provide links to support their stance and recommended news content related to the opinion. Group 3 identified the same issue but considered an artificial intelligence agent (shown as a dog in Figure 4(a)) to suggest to the user that they could provide justifications. Their system also provided users with controls for who could view or react to their content. Interestingly, Group 3 introduced a character limit, focusing on shorter messages and propagation.
@@
[…]
A sketch showing a content posting page with separated boxes for presenting ones’ opinion, reasoning and supportive material.
Beyond providing written content and comments, two of the artefacts included a mechanism to react to the content. Group 1’s artefact used stars to show support, and such reactions would increase the topic’s visibility on lists, stating that users would be instructed to:
Presenting participant: [Describing user guidelines for researchers] Like topics which you find important so that they rise to the discussion [further explaining the sentence to researchers] so it uses the Reddit approach [emphasis ours]
As they explained, there was competition for visibility similar to the Reddit platform; where a highly upvoted content gains more visibility. Group 3 provided various reaction icons and the ability to recirculate content on their artefacts, which they explained to be used to indicate some kind of social support towards the posts’ author:
Researcher: Do I see correctly, that you have some kind of like buttons there?
Presenting participant: Yes
Researcher: Why did you want to add the like buttons?
Presenting participant: We wanted to ensure this [platform] had a positive atmosphere (laugher) and we wanted to make it so that you could show your support to others using these reactions [emphasis ours]
This articulation is arguably different from Group 1’s, as Group 3 did not consider reactions a mechanism for gaining popularity. Rather, they emphasised their meaning for socio-emotional support in the final presentation. In fact, they chose the reactions in a manner that they should not be able to be appropriated for displaying negative attitudes. Groups 2 and 4 did not introduce this kind of feature to the artefacts, with Group 2 articulating how reactions, such as ‘likes‘ are not beneficial for constructive discussion and lead to attention seeking, explaining in the final presentation that:
Presenting participant: But it’s precisely the idea that people questioned the significance of having someone who gets like 3,000 likes and that someone gets no likes, that people would focus on the discussion and not how it’s said, dopamine spike.
Recent academic discussions on social media have focused on content moderation and its challenges  (Rantasila et al., 2022). Oonly groups 1 and 3 reported any form of content moderation in the final artefact, even though the focus was to devise an online space for political discussion. Group 1 hints at rules for the platform, which are enforced by moderators selected by some authority, which exemplifies the intention to govern the content (Rantasila et al., 2022). While it remained unclear who is part of the moderation team, the moderation must be conducted as a team effort. Group 3 described a more user-driven approach to moderation with options to report posts, add trigger warnings, and show reminders to enforce rules and block users. Similarly, Group 3 mentioned that there is a human editor to monitor automated fact-checking. This shows a mix of hard (e.g., removal of reported posts) and soft (e.g. trigger warnings) moderation styles (Rantasila et al., 2022). While we can surmise rules against spam and hate speech from Group 3 suggesting these as features, no group clearly described what content is warranted or sanctioned on their sites.
@@ 4.3. Profiles
All four designs contained the idea of a persistent user identity as a core part of the platform. At the same time, there was a lot of diversity in the details that should be included in the artefacts. All sites included some profile pictures. Group 2 allowed pseudonyms and a person to change their username, which would keep the posting history intact. Instead, groups 1, 3 and 4 enforced a real-name policy. However, Group 3 hinted at the possibility of parody accounts, although marked by an indicator. The policy of enforcing clear names is often debated also on social media services such as Facebook (Haimson and Hoffmann, 2016).
User verification was discussed in all final presentations. Group 1 argued that the system should have strong authentication during the final presentation to support the use of the data and build up a stronger accountability for participants’ actions:
Researcher: Can I ask at this point why you ended up with strong authentication?
Presenting participant: Because we have a social media platform that is used as an element in official decision-making. [- -]
Presenting participant: Of course, to the extent to which strong authentication increases this status in the minds of many. It ensures that it cannot be a random caller to a gossip magazine, or a bully or so on. [The authentication] controls for this. This is where we ended up in this case.
but it also was an approach to ensure proper behaviour on the platform, and Group 3 argued in a corresponding manner that there should be a mechanism to ensure the poster ‘is a real person’ and not pretend to be some other person. Similar arguments were also presented by Group 2 and Group 4.
Interestingly, the differences do not stop here (see Table 3). There was variance in what aspects of users’ identities were made relevant in the profiles, spanning from elements such as free self-descriptions to pre-given political affiliations and even to proofs of deviating from platform’s rules. Profiles included a free-form description field in groups 3 and 4, allowing for freedom on identity statements. Group 2 did not design such an element; instead, after the user name, an indication of the user’s party affiliation was shown. Group 4 added age and occupation in addition to the free-formed field, as well as a list of users’ political interests. Group 3 instead included, for example, the number of warnings the user had received for breaking any rules.
@@ 4.4. From Features to Affordances to Ideologies
The following section describes the differences and similarities of the artefacts along four key affordances (each row in Table 4). In the following paragraphs, we will go through each key affordance and the differences and commonalities between groups, along with their ideological leanings.
@@
Popularity-based Visibility Encourage Refuse Discourage Not addressed
Open Network Interactions Refuse Refuse Encourage Refuse
User Intervention Allow Not addressed Encourage Encourage
@@
The groups varied in their stance towards the mechanics of content visibility on their platforms: the right-leaning group preferred popularity-based metrics to promote content on the platform. Posts would have a like count, prioritising their visibility on the platform and thus introducing a competitive logic for attention: each post would compete with others in gaining reactions. Therefore, the group encouraged popularity-based visibility by fostering a usage pattern that aims for posts that receive more reactions and consequently get rewarded with more visibility, similar to platforms like Reddit and Facebook (Van Raemdonck and Pierson, 2021). Both left-leaning Groups 2 and 3 rejected popularity-based design approaches. As we observed above, while Group 3 mentions the need for reactions and likes, they highlight them as easy ways to show support, not as a popularity-based logic. Therefore, Group 3 discouraged popularity-based visibility, but still had some affordances which may make it possible, whereas Group 2 outright refused it by not offering any affordance for users to engage with content based on its popularity.
@@
Regarding interaction forms, all but Group 3 opted for a subforum architecture. Group 1, 2 and 4 all decided to section the discussion on their platform according to organisational levels (Group 1); parties, society and others (Group 2); and themes or location (Group 4). While a platform technically allows anyone to see and interact with every user’s content, there is still a clear difference between a system requesting users to post in particular subforums and one accumulating posts on a news feed. The former can be considered closed group interactions because they happen in clearly delineated spaces, and content does not automatically reach users across a platform (Van Raemdonck and Pierson, 2021). Such architecture entrusts the site administrators and moderators responsible for sanctioning or rewarding certain behaviours, thus refusing open network interaction. Contrary to that, the affordances of Group 3’s artefact encourage open network interactions, as content moves freely across the platform. Such an architecture focuses on the sender instead of an overarching group space like a subforum. This puts more responsibility and power onto the senders, as they have more possibilities for intervention and structuring the debate (Van Raemdonck and Pierson, 2021). This shows a contrast to the other left-leaning Group 2, as they chose the older idea of forums as a solution for their platform architecture.
@@
Regarding user interventions, we observed various approaches and remarkable differences in the way the groups’ artefact afforded users to be involved. The right-leaning Group 1 considered the moderation rules so important that one of their artefacts was a list of written rules. While their system imposed a governing style to moderation (Rantasila et al., 2022), they still allowed user interventions through the option to report to content and initiatives for new subforums but did not encourage such action, which shows a more neutral stance of allowing. Instead, the left-leaning Group 3 used self-guided moderation strategies with trigger warnings and warning counts (which could lead to banning from the platform) (Rantasila et al., 2022). Further, together with the mixed Group 4, they thought of structured content input that promotes user in adding justification for their posts. Thus, both groups’ artefacts encourage user intervention.
@@
Finally, it seems that groups including left-leaning participants provided more flexibility for users to present themselves, encouraging or at least not restricting flexible profile work (DeVito et al., 2017). While all groups required a persistent identity from their users, features such as a changeable username (Group 2) or parody accounts (Group 3) offered users agency to express themselves more freely. Additionally, Group 3 offered users control over the visibility of their content, encouraging users to perform self-presentation by catering to different audiences (DeVito et al., 2017). In stark contrast, the right-leaning group offered the least opportunities to customise profiles or present freely, which indicates an expectation for the users to present persistent roles on their platform and an affordance that discourages flexible profile work.
@@
Our study only had one right-leaning and one mixed composition group due to participant recruitment challenges (see foonote 1). However, two left-leaning groups already allow us to highlight diversity within the same end of the political spectrum, demonstrating how various other factors may also impact design work. For example, with profile work, Group 3 was much more creative in terms of attributes (see Table 3). Further, Group 2 subscribes to the forum approach, which separates interactions on the platforms into pre-defined groups and gives less intervention capabilities to individual users.